OpenStudy (anonymous):

What is 2∕15ths ÷ 2∕5ths

OpenStudy (amistre64):

2∕15 ÷ 2∕5 = 2∕15 * (2∕5)^-1 = 2/15 * 5/2

myininaya (myininaya):

what is ths?

OpenStudy (anonymous):

So what is the answer

OpenStudy (amistre64):

the answer is what you get when you solve for 2/15 * 5/2

OpenStudy (amistre64):

i assume the division is what threw you and you needed a little help to get to a simpler form

OpenStudy (anonymous):

o ty

OpenStudy (amistre64):

yw, and good luck :)

OpenStudy (anonymous):